I do not understand how to solve
solve for x

x-a=-x+a+4b

Answer by LilSkittleMd(119)

x-a=-x+a+4b
First you need to get x on one side of the equation. Do this by adding x to both sides of the equation. This gives you:
2x-a=a+4b
Now add a to both sides of the equation. This gives you:
2x=2a+4b
Divide each term by 2 to isolate x. This will give you:
x=a+2b